OpenStudy (anonymous):

f(x)= ((6x)^−1/2) + 5 what is f ' (x)? if you use the power rule, please help me with the negative exponent, because I dont understand it

OpenStudy (rogue):

The power rule is as follows\[\frac {d}{dx} x^n = n x^{n-1}\]

OpenStudy (rogue):

Is your function\[f(x) = 6x^{-.5} + 5\]or\[f(x) = (6x)^{-.5} + 5\]?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

so it should be -.5(6x)^-3/2 should it not?

OpenStudy (rogue):

If its the second one, then its not a variable to a power, but rather a function to a power. So the formula you must follow requires use of the chain rule, as well as the power rule.\[\frac {d}{dx} u^n = n u^{n-1} \frac {du}{dx}\]

OpenStudy (anonymous):

so what is the answer then?

OpenStudy (rogue):

-3(6x)^(-3/2) I think

OpenStudy (rogue):

You just have to multiply your answer that you have from the power rule by the derivative of u (6x), which is just 6.

OpenStudy (rogue):

Actually, no, never mind, its still a variable to a power.

OpenStudy (rogue):

Use the normal power rule.

OpenStudy (rogue):

\[f(x) = \frac {x^{-.5}}{\sqrt {6}} + 5\]

OpenStudy (rogue):

Gah, annoying powers...

OpenStudy (rogue):

\[f'(x) = \frac {-1}{2} \frac {1}{\sqrt {6}} x^{-1.5} = \frac {-1}{2 \sqrt {6} x^{1.5}}\]

OpenStudy (anonymous):

try it the other way, it didnt wordk in the homework

OpenStudy (rogue):

That last answer I gave should have been correct, unless its not the right problem... :( http://www.wolframalpha.com/input/?i=d%2Fdx+%286x%29^%28-.5%29+%2B+5

OpenStudy (anonymous):

the answer was 1/9

OpenStudy (rogue):

Did they ask for f'(x) at a certain point?

OpenStudy (anonymous):

nevermind that was a different problem, but the other was still wrong kill me now!